paperheight  29.69577cm 845.04684pt \paperheight
paperwidth  20.99701cm 597.50787pt \paperwidth
printpaperheight  29.69577cm 845.04684pt \printpaperheight
printpaperwidth  20.99701cm 597.50787pt \printpaperwidth
topspace  2.4994cm  71.12546pt \topspace
backspace  2.49966cm 71.13275pt \backspace
height  24.99644cm 711.3191pt \makeupheight
width  14.99773cm 426.78743pt \makeupwidth
top  0.0cm  0.0pt  \topheight
topdistance  0.0cm  0.0pt  \topdistance
header  1.99962cm 56.90294pt \headerheight
headerdistance 0.35141cm 0.0pt \headerdistance
textheight  20.29437cm 577.51323pt \textheight
footerdistance 0.35141cm 0.0pt \footerdistance
footer  1.99962cm 56.90294pt \footerheight
bottomdistance  0.0cm  0.0pt  \bottomdistance
bottom  0.0cm  0.0pt  \bottomheight
leftedge  0.0cm  0.0pt  \leftedgewidth
leftedgedistance  0.0cm  0.0pt  \leftedgedistance
leftmargin  2.65602cm 75.58197pt \leftmarginwidth
leftmargindistance 0.42163cm 11.99829pt \leftmargindistance
textwidth  14.99773cm 426.78743pt \textwidth
rightmargindistance 0.42163cm 11.99829pt \rightmargindistance
rightmargin  2.65602cm 75.58197pt \rightmarginwidth
rightedgedistance  0.0cm  0.0pt  \rightedgedistance
rightedge  0.0cm  0.0pt  \rightedgewidth

bodyfontsize  12.0pt  \globalbodyfontsize
line  2.8ex  \normallineheight
height  .72  \strutheightfactor
depth  .28  \strutdepthfactor
topskip  1.0  \topskipfactor
maxdepth  0.4  \maxdepthfactor

1  1_{1  1}

paperheight  29.69577cm 845.04684pt \paperheight
paperwidth  20.99701cm 597.50787pt \paperwidth
printpaperheight  29.69577cm 845.04684pt \printpaperheight
printpaperwidth  20.99701cm 597.50787pt \printpaperwidth
topspace  2.4994cm  71.12546pt \topspace
backspace  2.49966cm 71.13275pt \backspace
height  24.99644cm 711.3191pt \makeupheight
width  14.99773cm 426.78743pt \makeupwidth
top  0.0cm  0.0pt  \topheight
topdistance  0.0cm  0.0pt  \topdistance
header  1.99962cm 56.90294pt \headerheight
headerdistance 0.35141cm 0.0pt \headerdistance
textheight  20.29437cm 577.51323pt \textheight
footerdistance 0.35141cm 0.0pt \footerdistance
footer  1.99962cm 56.90294pt \footerheight
bottomdistance  0.0cm  0.0pt  \bottomdistance
bottom  0.0cm  0.0pt  \bottomheight
leftedge  0.0cm  0.0pt  \leftedgewidth
leftedgedistance  0.0cm  0.0pt  \leftedgedistance
leftmargin  2.65602cm 75.58197pt \leftmarginwidth
leftmargindistance 0.42163cm 11.99829pt \leftmargindistance
textwidth  14.99773cm 426.78743pt \textwidth
rightmargindistance 0.42163cm 11.99829pt \rightmargindistance
rightmargin  2.65602cm 75.58197pt \rightmarginwidth
rightedgedistance  0.0cm  0.0pt  \rightedgedistance
rightedge  0.0cm  0.0pt  \rightedgewidth

bodyfontsize  12.0pt  \globalbodyfontsize
line  2.8ex  \normallineheight
height  .72  \strutheightfactor
depth  .28  \strutdepthfactor
topskip  1.0  \topskipfactor
maxdepth  0.4  \maxdepthfactor

2  2_{2  2}

paperheight  29.69577cm 845.04684pt \paperheight
paperwidth  20.99701cm 597.50787pt \paperwidth
printpaperheight  29.69577cm 845.04684pt \printpaperheight
printpaperwidth  20.99701cm 597.50787pt \printpaperwidth
topspace  2.4994cm  71.12546pt \topspace
backspace  2.49966cm 71.13275pt \backspace
height  24.99644cm 711.3191pt \makeupheight
width  14.99773cm 426.78743pt \makeupwidth
top  0.0cm  0.0pt  \topheight
topdistance  0.0cm  0.0pt  \topdistance
header  1.99962cm 56.90294pt \headerheight
headerdistance 0.35141cm 0.0pt \headerdistance
textheight  20.29437cm 577.51323pt \textheight
footerdistance 0.35141cm 0.0pt \footerdistance
footer  1.99962cm 56.90294pt \footerheight
bottomdistance  0.0cm  0.0pt  \bottomdistance
bottom  0.0cm  0.0pt  \bottomheight
leftedge  0.0cm  0.0pt  \leftedgewidth
leftedgedistance  0.0cm  0.0pt  \leftedgedistance
leftmargin  2.65602cm 75.58197pt \leftmarginwidth
leftmargindistance 0.42163cm 11.99829pt \leftmargindistance
textwidth  14.99773cm 426.78743pt \textwidth
rightmargindistance 0.42163cm 11.99829pt \rightmargindistance
rightmargin  2.65602cm 75.58197pt \rightmarginwidth
rightedgedistance  0.0cm  0.0pt  \rightedgedistance
rightedge  0.0cm  0.0pt  \rightedgewidth

bodyfontsize  12.0pt  \globalbodyfontsize
line  2.8ex  \normallineheight
height  .72  \strutheightfactor
depth  .28  \strutdepthfactor
topskip  1.0  \topskipfactor
maxdepth  0.4  \maxdepthfactor

3  3_{3  3}

paperheight  29.69577cm 845.04684pt \paperheight
paperwidth  20.99701cm 597.50787pt \paperwidth
printpaperheight  29.69577cm 845.04684pt \printpaperheight
printpaperwidth  20.99701cm 597.50787pt \printpaperwidth
topspace  2.4994cm  71.12546pt \topspace
backspace  2.49966cm 71.13275pt \backspace
height  24.99644cm 711.3191pt \makeupheight
width  14.99773cm 426.78743pt \makeupwidth
top  0.0cm  0.0pt  \topheight
topdistance  0.0cm  0.0pt  \topdistance
header  1.99962cm 56.90294pt \headerheight
headerdistance 0.35141cm 0.0pt \headerdistance
textheight  20.29437cm 577.51323pt \textheight
footerdistance 0.35141cm 0.0pt \footerdistance
footer  1.99962cm 56.90294pt \footerheight
bottomdistance  0.0cm  0.0pt  \bottomdistance
bottom  0.0cm  0.0pt  \bottomheight
leftedge  0.0cm  0.0pt  \leftedgewidth
leftedgedistance  0.0cm  0.0pt  \leftedgedistance
leftmargin  2.65602cm 75.58197pt \leftmarginwidth
leftmargindistance 0.42163cm 11.99829pt \leftmargindistance
textwidth  14.99773cm 426.78743pt \textwidth
rightmargindistance 0.42163cm 11.99829pt \rightmargindistance
rightmargin  2.65602cm 75.58197pt \rightmarginwidth
rightedgedistance  0.0cm  0.0pt  \rightedgedistance
rightedge  0.0cm  0.0pt  \rightedgewidth

bodyfontsize  12.0pt  \globalbodyfontsize
line  2.8ex  \normallineheight
height  .72  \strutheightfactor
depth  .28  \strutdepthfactor
topskip  1.0  \topskipfactor
maxdepth  0.4  \maxdepthfactor

4  4_{4  4}

Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country.

Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country.

Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country.

Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country.

Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country. Now is the time for all good men to come to the aid of their country.